Circular No. 6607

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) L. Woodney, University of Maryland (UM); J. McMullin, National

Radio Astronomy Observatory (NRAO); M. A’Hearn, UM; and N. Samarasinha, National Optical Astronomy Observatory, write: “In observations at the NRAO 12-m telescope on Mar. 23 UT, we detected H_{2}CS for the first time in C/1995 O1. The 7(16)–6(15) line at 244.0478 GHz had a peak (corrected main-beam brightness) temperature of 0.02 K, and FWHM of 3 km/s. Assuming a rotational temperature of 80 K, we derive a production rate of 5.4 × 10^{26}. OCS J(12–11) measured on Mar. 22 had a production rate of 6.1 × 10^{27} mol/s.”

M. K. Bird and P. Janardhan, University of Bonn; and P. Gemsheimer, W. Huchtmeier and T. L. Wilson, Max-Planck-Institut für Radioastronomie (MPIfR), Bonn, report detection of ammonia in comet C/1995 O1 using the 100-m MPIfR radio telescope at Effelsberg. The following peak mainbeam brightness temperatures were detected for three individual K-band lines of NH_{3}: (3,3) at 23.870 GHz, 0.42 K; (2,2) at 23.723 GHz, 0.26 K; (1,1) at 23.695 GHz, 0.13 K. The standard error for all three measurements is 0.04 K. A main-beam brightness temperature of 1 K is equivalent to a flux density of 0.7 Jy for observations at these frequencies. The lines are centered symmetrically on the velocity of the cometary nucleus. The full width at half-maximum of the (3,3) line is 1.6 km/s. No noticeable changes in the strengths of the lines could be detected between the two observation dates on Mar. 13.4 and 25.5.

1991 VH P. Pravec, M. Wolf, and L. Šarounová, Ondřejov Observatory, report:

“Our photometric observations of this Apollo object on ten nights during Feb. 27–Mar. 13 revealed the presence of two periods in the lightcurve, 2.6236 ± 0.0003 and 32.74 ± 0.05 hr. Analysis of the lightcurve suggests that, like 1994 AW_{1} (which has a rather similar orbit), 1991 VH may be an eclipsing/occulting binary (Pravec and Hahn 1997, Icarus, in press). Further observations of the eclipse/occultation events from more than one station are desirable. A current ephemeris for the object is on MPC 27570. Approximate ephemerides for the primary and secondary minima are JD 2450518.38 + 1.364E and 2450520.37 + 1.364E, respectively.

See http://sunkl.asu.cas.cz/~ppravec/91vh.html for details.”

1997 March 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green